### v2

# Langevin equation with potential of mean force: The case of anchored bath

Abstract
The potential of mean force (PMF) is an effective average potential acting on an open system, renormalized due to the interaction with the surrounding thermal bath. The PMF determines the correction to the equilibrium Gibbs distribution, but it is generally unclear how to implement the concept for the time-dependent phenomena described by a (generalized) Langevin equation. We study a model where the system is a single particle (so there are no complications related to internal forces) and a non-trivial PMF is due to the presence of on-site (anchor) potentials applied to the bath particles. We found that the PMF does not merely replace the external potential, but also makes the dissipation kernel and statistical properties of noise dependent on the system’s position. That dependence is determined by the internal bath and system-bath interactions and is a priori unknown. Therefore, in the general case the Langevin equation with the PMF is not closed and thus inoperable. However, for systems with linear forces the aforementioned dependence on the system’s position is canceled. As an example, we consider a model where the bath is formed by the Klein-Gordon chain, i. e. a harmonic chain with on-site harmonic potentials. In that case, the generalized Langevin equation has the standard form with an external potential replaced by a quadratic PMF.
I Introduction
When addressing the dynamics of open systems, one often assumes that the environment, or the thermal bath, is passive in the sense that it does not produce any systematic force on the system kept at rest. A passive bath is still responsible for the velocity-dependent dissipative force on the moving system, but the only force it exerts on the fixed system is Langevin noise that vanishes after an appropriate averaging. The passivity of a thermal bath does not necessarily imply that the system-bath interaction is weak, but often holds due to symmetry. The required symmetry is not a property of the bath alone, but of the bath and system combined. For example, in conceptual models of Brownian motion the system is a single point-like particle, the bath is a homogeneous fluid, and a potential describing the system-bath interaction is pairwise. For that setting, the average bath-induce force on a fixed system vanishes because contributions from different bath molecules are canceled out. As a result, the bath is passive for an arbitrary strength of the system-bath interaction.
Now, in the above example, let us replace the point-like particle by a dumbbell-like system consisting of two particles, say $1$ and $2$, and consider the force the bath acts on particle $1$. That force generally does not vanish after averaging because particle $1$ interacts with the bath whose symmetry is broken by the presence of particle $2$. Generalizing, for any system consisting of more than one particle, the assumption of the passive bath does not generally hold because one part of the system breaks the symmetry of the bath interacting with other parts of the system. Under certain conditions, the effect may be negligible, but for many systems studied in chemical physics it is often essential Kirkwood ; DO ; Roux .
In this paper, we consider a model in which molecules of the bath experience, in addition to pairwise potentials from other molecules, also individual on-site potentials. That potentials, tending to localize bath molecules, break translational invariance of the bath. As a result, the bath is non-passive even when the system is a single particle.
When the bath is passive, the equilibrium distribution for the system has the canonical form
$\rho_{s}=\frac{1}{Z_{s}}\,e^{-\beta H_{s}},\quad Z_{s}=\operatorname{Tr}_{s}\left(e^{-\beta H_{s}}\right),$ (1)
where $H_{s}$ is the bare Hamiltonian of the system and $\operatorname{Tr}_{s}(\cdots)$ denotes integration over the sets of the system’s coordinates $X$ and momenta $P$ (we assume that the system is classical). For a non-passive bath $\rho_{s}$ involves the re-normalized Hamiltonian of the system $H_{s}^{*}$, called the Hamiltonian of mean force (HMF)  Kirkwood ; DO ; Roux ; Gelin ; Lutz ; Talkner
$\rho_{s}=\frac{1}{Z_{s}^{*}}\,e^{-\beta H_{s}^{*}},\quad Z_{s}^{*}=\operatorname{Tr}_{s}\left(e^{-\beta H_{s}^{*}}\right).$ (2)
In classical models, the renormalization affects only the system’s potential energy $V(X)$, adding to it a bath-induced correction $\Delta V$. Then the HMF can be written as
$H_{s}^{*}=K(P)+V_{*}(X),$ (3)
where $K(P)$ is the system’s kinetic energy and $V_{*}(X)$ is the renormalized potential
$V_{*}(X)=V(X)+\Delta V(X),$ (4)
called the potential of mean force (PMF). The bath-induced potential $\Delta V$ in general depends on temperature and quantifies the deviation of the equilibrium distribution from the canonical form (1).
As a note on nomenclature, the term PMF is sometimes reserved to $\Delta V$, or even to the total renormalized Hamiltonian of the system $H_{s}^{*}$ (thus identifying the PMF and the HMF). However, these definitions do not appear to be very common. Our definition of the PMF as a total effective potential acting on the system, Eq. (4), is the same as in Ref. Talkner . A drawback of such terminology is that it does not offer a generally-accepted term for the key quantity $\Delta V$. We shall refer to $\Delta V$ as the bath-induced potential and call the bath non-passive if it generates a non-zero (and non-constant) $\Delta V$. We avoid the term “active bath” which may suggest irrelevant connotations with the realm of active matter, which is beyond the scope of the paper.
As discussed in the next section, the PMF is defined as a quantity that characterizes the equilibrium distribution and determines average equilibrium properties of the system. Clearly, the PMF must also affect time-dependent properties, such as time correlation and response functions, described by a Langevin equation - but how it affects them is not always obvious. The simplest assumption is that the Langevin equation and the corresponding fluctuation-dissipation relation retain their standard form and the only required modification is the replacement of the bare system potential by the PMF. That approach, broadly adopted in empirical studies, received theoretical endorsement in Refs. Lange ; Kinjo , where the standard generalized Langevin equation with the PMF was derived from the underlying dynamics using the projection operator technique. However, those derivations were later criticized in Refs. Schilling ; Schilling_review , with the conclusion that for systems with a PMF the standard Langevin equation can be derived neither exactly nor using any well-controlled approximation. On the other hand, in Refs. Vroy ; Vroy2 ; Lee it was argued that a Langevin equation actually can be recovered from the first principles, but only at the expense of the noise and the dissipation kernel being position-dependent and with a modified fluctuation-dissipation relation. That conclusion seems to corroborate the earlier results for a system of several Brownian particles DO . Using a different approach, a generalized Langevin equation with an additional dissipation force, independent of the system’s velocity, was suggested in Netz .
In all studies cited above, the derivation of the Langevin equation with a PMF was addressed for many-particle systems with internal forces, which significantly complicates the analysis. In this paper we consider a model where the system is a single Brownian particle with no internal degrees of freedom. In that case, as we mentioned above, a translationally-invariant bath is passive due to symmetry, and thus the PMF is redundant. In our model, the translational invariance of the bath is broken by on-site (or anchor) potentials applied to the particles of the bath. Such a bath is not passive; it exerts a systematic average force on a fixed system. Our goal is to determine how this force affects the Langevin equation for the system.
Our interest in such a model is threefold. First, the model involves a non-trivial PMF but has no complications related to internal forces in the system. Second, the Langevin dynamics with an anchored bath is relevant for a variety of physical systems, see Savin ; Flach ; Prem ; Dhar ; Flach2 and references therein. For example, in complex molecular structures like proteins, the lighter and heavier units may be interpreted as the bath degrees of freedom and anchor (substrate) centers, respectively Kneller ; Xie ; Milster . Third, the spectrum of an anchored bath has both upper and lower bounds, which may lead to peculiar ergodic properties of the Langevin dynamics Plyukhin_NEO .
We shall consider the model first for the general and then for a more specific settings. For the general setting, we make no assumptions about specific forms of the potentials involved. For that general case, we are unable to derive a generalized Langevin equation in the standard form. The equation we derived instead involves noise and the dissipation kernel, which depend on the system’s position in a way which is a priori unknown. As a result, the equation is not closed and practically unmanageable.
On the other hand, the situation is simplified significantly for a specific setting when the bath is formed by the harmonic Klein-Gordon chain. Due to the linearity of that model, the aforementioned dependence of the dissipation kernel on the system’s position is canceled, and the generalized Langevin equation has the standard form with a quadratic PMF.
The layout of the paper is as follows. In Section 2 we review the PMF. Although the concept is quite mature, there is some ambiguity and variations in the literature, so we feel a brief review might be useful. In Section 3 we outline the model with a single-particle system and anchored bath. In Sections 4 and 5 we attempt to derive the Langevin equation for the general setting using the Mazur-Oppenheim projection operator technique. In Section 6 we describe a more specific setting with the Klein-Gordon lattice as a bath and evaluate the corresponding PMF. In Section 7 we show that for the Klein-Gordon model the Langevin equation derived in Section 5 is reduced to the standard form. In Appendix we derived that equation by direct integration of bath variables.
II Review of PMF
Consider a classical system, with bare Hamiltonian $H_{s}(X,P)$, interacting with a thermal bath with bare Hamiltonian $H_{b}(x,p)$ via potential $\Phi(X,x)$. Here $(X,P)$ and $(x,p)$ denote the sets of coordinates and momenta for the system and the bath, respectively. The total Hamiltonian $H=H_{s}+H_{b}+\Phi$ is convenient to compose as
$H=H_{s}+H_{0}(X),\quad H_{0}(X)=H_{b}+\Phi,$ (5)
where $H_{0}(X)$ is the Hamiltonian of the bath interacting with the system with fixed position vector $X$.
Suppose the overall system (the system and the bath combined) is in thermal equilibrium with canonical distribution
$\rho=\frac{1}{Z}\,e^{-\beta H}=\frac{1}{Z}\,e^{-\beta(H_{s}+H_{0})},\quad Z=\operatorname{Tr}_{sb}\left(e^{-\beta H}\right),$ (6)
where $\operatorname{Tr}_{sb}(\cdots)$ denotes integration over phase-space variables of both the system and the bath. We are interested in the reduced (marginal) distribution $\rho_{s}$ for the system only,
$\rho_{s}=\operatorname{Tr}_{b}(\rho)=\frac{Z_{0}(X)}{Z}\,e^{-\beta H_{s}}.$ (7)
Here $\operatorname{Tr}_{b}(\cdots)$ denotes integration over phase-space variables of the bath, and $Z_{0}(X)$ is the partition function for the canonical distribution $\rho_{0}$ corresponding to $H_{0}$,
$\rho_{0}=\frac{1}{Z_{0}(X)}\,e^{-\beta H_{0}},\quad Z_{0}(X)=\operatorname{Tr}_{b}\left(e^{-\beta H_{0}}\right).$ (8)
Distribution $\rho_{0}$ is the equilibrium distribution for the bath connected to the system fixed at position $X$, so that partition function $Z_{0}$ in general depends on $X$. Since that dependence is a priori unknown, expression (7) for $\rho_{s}$ is not directly useful or evocative. For some special settings the total partition function can be, exactly or approximately, factorized, $Z=Z_{s}\,Z_{0}$. In that case, $Z_{0}$ is canceled out in Eq. (7), and $\rho_{s}$ takes the Boltzmann-Gibbs form (1),
$\rho_{s}=\frac{1}{Z_{s}}\,e^{-\beta H_{s}},\quad Z_{s}=\operatorname{Tr}_{s}\left(e^{-\beta H_{s}}\right).$ (9)
However, in the general case the aforementioned factorization of $Z$ does not hold, and expression (7) for $\rho_{s}$ is not reduced to the form (9). Yet it is possible to present the general expression (7) in a more suggestive form as follows.
Integrating over the bath variables, the total partition function can be written as
$Z=\operatorname{Tr}_{sb}[e^{-\beta(H_{s}+H_{0})}]=\operatorname{Tr}_{s}[e^{-\beta H_{s}}\,Z_{0}(X)],$ (10)
then Eq. (7) takes the form
$\rho_{s}=\frac{Z_{0}(X)}{\operatorname{Tr}_{s}[e^{-\beta H_{s}}\,Z_{0}(X)]}\,e^{-\beta H_{s}}.{}$ (11)
It is convenient to parameterize the dependence of $Z_{0}$ on $X$ as
$Z_{0}(X)=c\,e^{-\beta\Delta V(X)}.$ (12)
While $Z_{0}$ has no direct physical meaning, the new function
$\Delta V(X)=-\frac{1}{\beta}\,\ln\left[\frac{Z_{0}(X)}{c}\right],$ (13)
can be interpreted as an effective bath-induced potential acting on the system in equilibrium. Indeed, substituting (12) into (11) gives
$\rho_{s}=\frac{1}{Z_{s}^{*}}\,e^{-\beta\,H_{s}^{*}},\quad Z_{s}^{*}=\operatorname{Tr}_{s}[e^{-\beta H_{s}^{*}}]$ (14)
where
$H_{s}^{*}=H_{s}+\Delta V(X)$ (15)
is called the Hamiltonian of mean force (HMF). By writing the bare Hamiltonian of the system as the sum of kinetic and potential energies $H_{s}=K(P)+V(X)$, the HMF can be presented as
$H_{s}^{*}=K(P)+V_{*}(X),\quad V_{*}(X)=V(X)+\Delta V(X),$ (16)
The renormalized potential $V_{*}(X)$ is commonly referred to as the potential of mean force (PMF). It is the sum of the potential energy $V$ of the bare system (which describes both the internal and external forces for the system disconnected from the bath) and the bath-induced potential $\Delta V$.
One may find slightly different definitions of the PMF in the literature. In particular, some authors identify the PMF with the bath-induced potential $\Delta V$ and do not include in the PMF the potential of the bare system $V$ DO . Our definitions of the HMF and the PMF by Eq. (16) coincide with those in Ref. Talkner . There are also differences in the literature about specific choices for parameter $c$ in Eq. (12); we shall discuss that below.
The interpretation of $\Delta V$ as the bath-induced effective potential can be further specified considering the force $F_{\alpha}$, exerted by the bath on $\alpha$-th particle of the system with the position vector $X_{\alpha}$,
$F_{\alpha}=-\frac{\partial H_{0}}{\partial X_{\alpha}}=-\frac{\partial\Phi}{\partial X_{\alpha}}.$ (17)
Consider the average force $\langle F_{\alpha}\rangle$, where the average is taken over bath variables with distribution $\rho_{0}$, Eq. (8),
$\langle F_{\alpha}\rangle=\operatorname{Tr}_{b}(\rho_{0}\,F_{\alpha})=\frac{1}{Z_{0}(X)}\,\int dx\,dp\,e^{-\beta H_{0}}\,F_{\alpha}.$ (18)
Differentiating Eq. (13), one finds
$\langle F_{\alpha}\rangle=-\frac{\partial}{\partial X_{\alpha}}\,\Delta V(X).$ (19)
Therefore, $\Delta V$ can be interpreted as a potential of the average force exerted by the bath on the fixed system. As was mentioned in the Introduction, for a one-particle system in a fluid that force vanishes by symmetry, but in the general case it is not zero.
Note that the total average force on the $\alpha$-th particle of the system, i. e. the sum of internal, external, and bath-induced forces,
$\langle F_{\alpha}^{tot}\rangle=-\left\langle\frac{\partial H}{\partial X_{\alpha}}\right\rangle=-\frac{\partial V}{\partial X_{\alpha}}+\langle F_{\alpha}\rangle,$ (20)
is given by the gradient of the PMF,
$\langle F_{\alpha}^{tot}\rangle=-\frac{\partial}{\partial X_{\alpha}}(V+\Delta V)=-\frac{\partial}{\partial X_{\alpha}}\,V_{*}(X).$ (21)
If the system is a single Brownian particle (as we assume in the sections to follow), the index $\alpha$ in the above equations is redundant and to be dropped.
We still need to specify the parameter $c$ in definition (13) of the bath-induced potential $\Delta V$. In that expression, $c$ determines only an additive constant (with respect to $X$) and, therefore, can be chosen arbitrarily. A common and often convenient choice is
$c=Z_{b},$ (22)
where $Z_{b}$ is the partition function for the equilibrium distribution of the isolated (not connected to the system) bath,
$\rho_{b}=\frac{1}{Z_{b}}\,e^{-\beta H_{b}},\quad Z_{b}=\operatorname{Tr}_{b}\left(e^{-\beta H_{b}}\right).$ (23)
In that case Eq. (13) gives for the bath-induced potential the following expression
$\Delta V=-\frac{1}{\beta}\ln\left[\frac{Z_{0}}{Z_{b}}\right]=-\frac{1}{\beta}\ln\left[\frac{\operatorname{Tr}_{b}\left(e^{-\beta H_{0}}\right)}{\operatorname{Tr}_{b}\left(e^{-\beta H_{b}}\right)}\right],$ (24)
and the HMF $H_{s}^{*}=H_{s}+\Delta V$ may be expressed as
$H_{s}^{*}=-\frac{1}{\beta}\ln\left[\frac{\operatorname{Tr}_{b}\left(e^{-\beta H}\right)}{\operatorname{Tr}_{b}\left(e^{-\beta H_{b}}\right)}\right],$ (25)
where $H=H_{s}+H_{0}$ is the total Hamiltonian of the combined system and the bath.
The advantage of the choice (22) can be seen considering the weak-coupling limit, when
$H_{0}=H_{b}+\Phi\to H_{b},\quad H=H_{s}+H_{0}\to H_{s}+H_{b},$ (26)
and therefore
$Z_{0}(X)\to Z_{b},\quad Z\to Z_{s}Z_{b}.$ (27)
In that limit the above equations give
$\Delta V\to 0,\quad V^{*}\to V,\quad H_{s}^{*}\to H_{s}.$ (28)
That is a plausible result: For a negligible system-bath interaction the bath-induced potential vanishes, and the equilibrium statistics of the system is determined by the system’s bare Hamiltonian $H_{s}$.
Let us note that expression (24) for $\Delta V$ often appears in the literature in a different form. Since
$\frac{Z_{0}}{Z_{b}}=\frac{1}{Z_{b}}\,\operatorname{Tr}_{b}\left(e^{-\beta\,H_{b}}\,e^{-\beta\Phi}\right)=\langle e^{-\beta\Phi}\rangle_{b},$ (29)
where the average is taken over bath variables with distribution $\rho_{b}$, $\langle\cdots\rangle_{b}=\operatorname{Tr}_{b}(\cdots\rho_{b})$, the expression (24) can be also written as
$\Delta V=-\frac{1}{\beta}\,\ln\langle e^{-\beta\Phi}\rangle_{b}.$ (30)
Similarly, the HMF (25) can be expressed as
$H_{s}^{*}=-\frac{1}{\beta}\,\ln\langle e^{-\beta(H_{s}+\Phi)}\rangle_{b}.$ (31)
Although Eqs. (30) and (31) may be convenient in some cases Talkner , they will not appear in the rest of the paper.
Had we chosen the parameter $c$ in a form different than (22), the bath-induced potential $\Delta V(X)$ would not vanish in the weak-coupling limit but instead reduce to a nonzero constant,
$\Delta V(X)=-\frac{1}{\beta}\ln\left[\frac{Z_{0}(X)}{c}\right]\to-\frac{1}{\beta}\ln\left[\frac{Z_{b}}{c}\right].$ (32)
That might be inconvenient, yet it is not incorrect, since observable physical quantities do not depend on $c$, provided $c$ is a constant with respect to $X$ and $P$. In Ref. Gelin , the choice $c=1$ was adopted. It was criticized in Talkner on the ground that for such a choice $\Delta V$ does not vanish in the weak coupling limit. In our view, the only defect of the choice $c=1$ is that the argument of the logarithm in Eq. (32) is not dimensionless.
In what follows, we shall consider a model for which $H_{0}$ has the special form
$H_{0}=h_{1}(x,X)+h_{2}(X),$ (33)
where $h_{2}(X)$ depends only on the system’s position $X$, while $h_{1}(x,X)$ depends on bath variables and $X$, yet the corresponding partition function
$Z_{1}=\operatorname{Tr}_{b}\left(e^{-\beta h_{1}}\right)$ (34)
does not depend on $X$. In that case,
$Z_{0}(X)=Z_{1}\,e^{-\beta\,h_{2}(X)},$ (35)
and it is convenient to define $c=Z_{1}$. Then, according to (13), the bath-induced potential is equal to $h_{2}(X)$,
$\Delta V(X)=-\frac{1}{\beta}\ln\left[\frac{Z_{0}(X)}{Z_{1}}\right]=h_{2}(X).$ (36)
III Anchored bath model
As discussed in the Introduction, the absence of a PMF in simplified models of Brownian motion is to some extent accidental: If the system of interest is a single particle with no internal degrees of freedom and the bath is a translationally invariant fluid, then the average force from the bath on the system is zero by symmetry. Our goal here is to study a model where the system of interest is still a structureless Brownian particle, but the translational invariance of the bath is broken, which results in a non-trivial PMF.
The bare Hamiltonian of the system is
$H_{s}=\frac{P^{2}}{2M}+V_{ex}(X),$ (37)
where $V_{ex}(X)$ is the external potential unrelated to the bath. The bath consists of a large number $N$ of particles and has the bare Hamiltonian
$H_{b}=\sum_{i=1}^{N}\frac{p_{i}^{2}}{2m}+\sum_{i>j}^{N}U(x_{i}-x_{j})+\sum_{i=1}^{N}U_{a}(x_{i}).$ (38)
Here, the first two sums represent the kinetic and pairwise potential energies. The key feature of the model is the presence of the third sum, which implies that each particle in the bath is subject to the additional “anchor” potential $U_{a}(x_{i})$. A particular realization of such a model, where the bath is an anchored harmonic chain, is sketched in Fig. 1. It is intuitively clear that the anchored bath is non-passive and generates a PMF which tends to localize the system. Physically, the model of the anchored bath may be relevant for complex heterogeneous systems, like polymer networks and gels, characterized by a broad hierarchy of relaxation times Milster . In such systems, heavier (slower) structural units may be essentially immobile on a relevant time scale and act as fixed anchoring sites for lighter and faster units.
We shall assume that the system-bath interaction energy $\Phi$ is pairwise,
$\Phi=\sum_{i=1}^{N}\phi(X-x_{i}),$ (39)
and write the total Hamiltonian in the form (5),
$H=H_{s}+H_{0},\quad H_{0}(X)=H_{b}+\Phi(X),$ (40)
where $H_{0}(X)$ is the Hamiltonian of the bath in the potential of the Brownian particle held fixed at position $X$.
The total Liouville operator for the model is convenient to write as
$L=L_{s}+L_{0}$ (41)
where $L_{s}$ and $L_{0}$ correspond to Hamiltonians $H_{s}$ and $H_{0}$, respectively,
$L_{s}=\frac{P}{M}\frac{\partial}{\partial X}+(F+F_{ex})\frac{\partial}{\partial P},\quad L_{0}=\sum_{i}\left\{\frac{p_{i}}{m}\frac{\partial}{\partial x_{i}}+f_{i}\frac{\partial}{\partial p_{i}}\right\}.$ (42)
Here $F$ and $F_{ex}$ are the forces on the particle due to the interaction with the bath and the external potential, respectively,
$F=-\frac{\partial\Phi}{\partial X},\quad F_{ex}=-\frac{\partial V_{ex}}{\partial X},$ (43)
and $f_{i}$ is the force on $i$-th particle of the bath
$f_{i}=-\frac{\partial H_{0}}{\partial x_{i}},$ (44)
which parametrically depends on the system’s position $X$.
We adopt the usual assumption of the Brownian motion theory that masses $m$ of bath particles are much smaller than the mass of the system, $m\ll M$. In that case, the bath is expected to evolve much faster than the system and quickly adjust to the instantaneous system’s position. Therefore, one may expect that the bath’s dynamics is close to that described by the Liouville operator $L_{0}$, which describes the bath in the field of a fixed (or, equivalently, an infinitely heavy) system. The operator $L_{0}$ has the properties
$L_{0}H_{0}=0,\quad L_{0}\,\varphi(H_{0})=0,$ (45)
where $\varphi(x)$ is an arbitrary differentiable function. A related property is that $L_{0}$ and $\varphi(H_{0})$ commute,
$L_{0}\,\varphi(H_{0})A=\varphi(H_{0})\,L_{0}A,$ (46)
where $A$ is an arbitrary dynamical variable. These properties of $L_{0}$ are generic and not affected by the presence of the anchor potential.
In the next two sections we shall use the Mazur-Oppenheim projection operator technique MO ; DO ; KO to derive a Langevin equation with the PMF for the model described in this section. The derivation is convenient to present in two steps. The first step is to re-write the equation of motion for the system separating the systematic and noisy components of the total force on the system. The result is what can be called the pre-Langevin equation, see Eq. (70) below. That equation is exact but impractical because it involves the system’s variables in an implicit form. The second step, discussed in Section 5, is to make that dependence explicit by applying an appropriate perturbation technique.
IV Pre-Langevin equation
The Mazur-Oppenhein technique MO is based on the projection operator $\mathcal{P}$ which averages an arbitrary dynamical variable $A$ over initial values of the bath variables
$\mathcal{P}A=\langle A\rangle=\operatorname{Tr}_{b}({\rho_{0}\,A})$ (47)
with distribution (8)
$\rho_{0}=\frac{1}{Z_{0}(X)}\,e^{-\beta H_{0}},\quad Z_{0}(X)=\operatorname{Tr}_{b}\left(e^{-\beta H_{0}}\right).$ (48)
Recall that $\rho_{0}$ is the equilibrium distribution of the bath in the field of the system fixed at $X$. This makes the projection (47) particularly relevant in the present context, because $\rho_{0}$ is the distribution used in the definition of the average force related to the PMF, see Eqs. (18)-(21).
The projection property $\mathcal{P}^{2}=\mathcal{P}$ is obvious, and another key property
$\mathcal{P}L_{0}=0$ (49)
follows from (46) with $\varphi(H_{0})=\rho_{0}$. The complementary projection operator
$\mathcal{Q}=1-\mathcal{P}$ (50)
has the properties
$\mathcal{P}\mathcal{Q}=\mathcal{Q}\mathcal{P}=0,\quad\mathcal{Q}L_{0}=L_{0}.$ (51)
The standard Mazur-Oppenheim approach assumes that the bath is passive, i. e.
$\mathcal{P}F=\langle F\rangle=0,$ (52)
where $F$ is the force exerted by the bath on the Brownian particle. Here we consider an extension when Eq. (52) does not hold.
The equation of motion for the system reads
$\dot{P}(t)=F(t)+F_{ex}(t),$ (53)
where $F$ is the force from the bath and $F_{ex}$ is the external force, see Eq. (43). The time dependence of forces can be described using the Liouville propagator
$F(t)=e^{Lt}\,F,\quad F_{ex}(t)=e^{Lt}\,F_{ex},$ (54)
where $F=F(0),F_{ex}=F_{ex}(0)$. We shall also use a similar notation $A=A(0)$ for the initial values of other dynamical variables. Inserting the identity operator $1=\mathcal{Q}+\mathcal{P}$ into expression (54) for $F(t)$,
$F(t)=e^{Lt}\mathcal{P}F+e^{Lt}\mathcal{Q}F=e^{Lt}\langle F\rangle+e^{Lt}\mathcal{Q}F$ (55)
brings the equation of motion (53) to the form
$\dot{P}(t)=e^{Lt}\left[\langle F\rangle+F_{ex}\right]+e^{Lt}\mathcal{Q}F.$ (56)
According to Eqs. (19) and (21), we can express $\langle F\rangle$ as the negative gradient of the the bath-induced potential $\Delta V$,
$\langle F\rangle=-\frac{\partial}{\partial X}\,\Delta V,$ (57)
and the total force $\langle F\rangle+F_{ex}$ as the negative gradient of the PMF $V_{*}(X)$,
$\langle F\rangle+F_{ex}=-\frac{\partial}{\partial X}\,V_{*}(X),\qquad V_{*}(X)=\Delta V(X)+V_{ex}(X).$ (58)
Then the equation of motion (56) can be written as
$\dot{P}(t)=-\nabla V_{*}[X(t)]+e^{Lt}\mathcal{Q}F,$ (59)
where the first term on the right-hand side is the gradient of the PMF at time $t$,
$\nabla V_{*}[X(t)]=e^{Lt}\frac{\partial}{\partial X}V_{*}(X)=\frac{\partial}{\partial X}V_{*}(X)\Big|_{X=X(t)}.$ (60)
That term is a conservative part of the total force acting on the system. The second term $e^{Lt}QF$ can be separated into a dissipative (non-conservative) force and noise. The usual way to facilitate such separation is to use the operator identity
$e^{(A+B)t}=e^{At}+\int_{0}^{t}d\tau\,e^{A(t-\tau)}B\,e^{(A+B)\tau}.$ (61)
With $A=L$ and $B=-\mathcal{P}L$, one splits the propagator into two parts,
$e^{Lt}=e^{\mathcal{Q}Lt}+\int_{0}^{t}d\tau\,e^{L(t-\tau)}\,\mathcal{P}L\,e^{\mathcal{Q}L\tau}.$ (62)
Substituting that into Eq. (59) and recalling that $\mathcal{P}L_{0}=0$ yields
$\dot{P}(t)=-\nabla V_{*}[X(t)]+F^{\dagger}(t)+\int_{0}^{t}d\tau\,e^{L(t-\tau)}\,\mathcal{P}L_{s}\,F^{\dagger}(\tau)$ (63)
where
$F^{\dagger}(t)=e^{\mathcal{Q}Lt}\mathcal{Q}F$ (64)
is often referred to as the projected force. Since $\mathcal{P}\mathcal{Q}=0$, the average $F^{\dagger}(t)$ is zero
$\langle F^{\dagger}(t)\rangle=\mathcal{P}F^{\dagger}(t)\sim\mathcal{PQ}=0,$ (65)
which suggests interpreting $F^{\dagger}(t)$ as noise. However, one has to keep in mind that $F^{\dagger}(t)$ depends on the system’s variables, and that dependence still needs to be revealed.
The integral term in Eq. (63) is expected to give rise to a dissipation force. Let us express $\mathcal{P}L_{s}F^{\dagger}(\tau)$ in the integrand in a more explicit form,
$\mathcal{P}L_{s}F^{\dagger}(\tau)=\frac{P}{M}\,\left\langle\frac{\partial}{\partial X}F^{\dagger}(\tau)\right\rangle+\frac{\partial}{\partial P}\,\langle F\,F^{\dagger}(\tau)\rangle.$ (66)
To evaluate the first term on the right-hand side, let us differentiate Eq. (65),
$\frac{\partial}{\partial X}\left\langle F^{\dagger}(\tau)\right\rangle=-\beta\left\langle\frac{\partial H_{0}}{\partial X}\,F^{\dagger}(\tau)\right\rangle-\langle F^{\dagger}(\tau)\rangle\frac{\partial}{\partial X}\ln Z_{0}+\left\langle\frac{\partial}{\partial X}F^{\dagger}(\tau)\right\rangle=0.$ (67)
Here we take into account that for a non-passive bath the distribution $\rho_{0}=Z_{0}^{-1}\exp(-\beta H_{0})$ depends on $X$ through both $H_{0}$ and $Z_{0}$. However, the term with $\frac{\partial}{\partial X}\ln Z_{0}$ vanishes because of the factor $\langle F^{\dagger}(\tau)\rangle=0$. As a result, one obtains from Eq. (67)
$\left\langle\frac{\partial}{\partial X}F^{\dagger}(\tau)\right\rangle=-\beta\left\langle F\,F^{\dagger}(\tau)\right\rangle,$ (68)
and Eq. (66) takes the form
$\mathcal{P}L_{s}F^{\dagger}(\tau)=\left(\frac{\partial}{\partial P}-\frac{\beta\,P}{M}\right)\langle F\,F^{\dagger}(\tau)\rangle.$ (69)
Substituting into (63) yields the equation of motion in the pre-Langevin form
$\dot{P}(t)=-\nabla V_{*}[X(t)]+F^{\dagger}(t)+\int_{0}^{t}d\tau\,e^{L(t-\tau)}\,\left(\frac{\partial}{\partial P}-\frac{\beta\,P}{M}\right)\langle F\,F^{\dagger}(\tau)\rangle.$ (70)
This equation is similar to that for a passive bath MO ; KO , with the only visible difference that the external potential $V_{ex}$ is replaced by the PMF $V_{*}=V_{ex}+\Delta V$. A hidden difference is that the projected force $F^{\dagger}(t)=e^{\mathcal{Q}Lt}\mathcal{Q}F$ has the additional factor $\mathcal{Q}$, which ensures that $\langle F^{\dagger}(t)\rangle\sim\mathcal{PQ}=0$ for any time. The equation is exact, but not yet operational because $F^{\dagger}(t)$ depends on the system’s variables. That dependence can be made explicit by expanding the projected force in powers of a small mass ratio parameter $\lambda$, $F^{\dagger}(t)=F_{0}(t)+\lambda\,F_{1}(t)+\cdots$. For a passive bath, in the leading (zeroth) order in $\lambda$ the projected force $F^{\dagger}(t)\approx F_{0}(t)$ does not depend on the system’s variables, and Eq. (70) is readily reduced to the standard generalized Langevin equation. However, as we shall see in the next section, for a non-passive bath $F^{\dagger}(t)$ depends on the system’s position even in the leading approximation. That dependence is a priori unknown, which may undermine the derivation of a practically useful Langevin equation with a PMF.
V Langevin equation
In order to bring the pre-Langevin Eq. (70) to a more amiable form, let us exploit a perturbation technique with the mass ratio
$\lambda^{2}=m/M$ (71)
as a small parameter. If the overall system is not too far from equilibrium, the momentum $P$ of the Brownian particle is close to the equilibrium value $\sqrt{M/\beta}$, and the typical momentum of bath particles is $\sqrt{m/\beta}$. Therefore, the scaled momentum $P_{*}=\lambda P$ of the Brownian particle is on average of the same order as momenta of the bath particles. To make the $\lambda$-dependence more explicit, let us re-write the pre-Langevin equation (70) in terms of the scaled momentum $P_{*}$,
$\dot{P}_{*}(t)=-\lambda\,\nabla V_{*}[X(t)]+\lambda\,F^{\dagger}(t)+\lambda^{2}\int_{0}^{t}d\tau\,e^{L(t-\tau)}\,\left(\frac{\partial}{\partial P_{*}}-\frac{\beta\,P_{*}}{m}\right)\langle F\,F^{\dagger}(\tau)\rangle.$ (72)
The Liouville operator (41) can be re-written as
$L=L_{0}+\lambda L_{1}$ (73)
where $L_{0}$ is still given by Eq. (42) and
$L_{1}=\frac{P_{*}}{m}\frac{\partial}{\partial X}+(F+F_{ex})\frac{\partial}{\partial P_{*}}.$ (74)
The projected force $F^{\dagger}(t)$ evolves as
$F^{\dagger}(t)=e^{\mathcal{Q}Lt}\mathcal{Q}F=e^{(L_{0}+\lambda\mathcal{Q}L_{1})t}\,\mathcal{Q}F,$ (75)
where we have used Eq. (51), $\mathcal{Q}L_{0}=L_{0}$. Using Eq. (61) with $A=L_{0}$ and $B=\lambda\mathcal{Q}L_{1}$ one finds
$e^{(L_{0}+\lambda\,\mathcal{Q}\,L_{1})\,t}=e^{L_{0}t}+\lambda\int_{0}^{t}d\tau e^{L_{0}\,(t-\tau)}\mathcal{Q}\,L_{1}\,e^{L_{0}\tau}+\mathcal{O}(\lambda^{2}).$ (76)
Substituting into (75) yields
$F^{\dagger}(t)=E_{0}(t)+\lambda E_{1}(t)+\mathcal{O}(\lambda^{2}),$ (77)
where
$E_{0}(t)=e^{L_{0}t}\mathcal{Q}F,\qquad E_{1}(t)=\frac{1}{m}\int_{0}^{t}d\tau\,P_{*}(t-\tau)\,e^{L_{0}(t-\tau)}\mathcal{Q}\frac{\partial}{\partial X}\,E_{0}(\tau),$ (78)
with both terms being zero-centered,
$\langle E_{0}(t)\rangle=\langle E_{1}(t)\rangle\sim\mathcal{PQ}=0.$ (79)
To order $\lambda^{2}$ the integral term in Eq. (72) is simplified and the equation takes the form
$\dot{P}_{*}(t)=-\lambda\,\nabla V_{*}[X(t)]+\lambda\,F^{\dagger}(t)-\frac{\lambda^{2}\beta}{m}\int_{0}^{t}d\tau\,P_{*}(t-\tau)\,\,e^{L(t-\tau)}\langle F\,E_{0}(\tau)\rangle.$ (80)
Since this equation is an approximation to order $\lambda^{2}$, it can only be used to evaluate properties which are either linear or quadratic in noise $\lambda\,F^{\dagger}(t)$. Upon averaging, the contributions linear in noise vanish, while contributions quadratic in noise are determined to order $\lambda^{2}$ by the leading term $E_{0}(t)$ in the expansion (77). Therefore, perturbation consistency requires to replace in Eq. (80) the projected force $F^{\dagger}(t)$ by the leading-order approximation $E_{0}(t)$. After that replacement, and after returning from the scaled to the true momentum $P=P_{*}/\lambda$, Eq. (81) takes the form
$\dot{P}(t)=-\nabla V_{*}[X(t)]+E_{0}(t)-\frac{\beta}{M}\int_{0}^{t}d\tau\,P(t-\tau)\,\,e^{L(t-\tau)}\langle F\,E_{0}(\tau)\rangle.$ (81)
The fluctuating force $E_{0}(t)$ can be written as
$E_{0}(t)=F_{0}(t)-\langle F\rangle,\quad F_{0}(t)=e^{L_{0}t}F$ (82)
and interpreted as the deviation of the force $F_{0}(t)$, exerted by the bath on the system fixed at $X$, from the average value $\langle F(X)\rangle$. Clearly, $E_{0}(t)$ does not depend on $P$, but it does depend on the system’s position $X$.
That dependence does not complicate the matter if the bath is a homogeneous non-anchored fluid. In that case, the dependence of forces on positions occurs only through the differences $X-x_{i}$ and $x_{i}-x_{j}$. When taking average over the bath coordinates $x_{i}$, one can change the integration variables $x_{i}^{\prime}=x_{i}-X$, so that the integrands depend on $x_{i}^{\prime}$ and $x_{i}^{\prime}-x_{j}^{\prime}$, but not on $X$. Therefore, although $F$ and $F_{0}(t)$ depend on $X$, their moments and correlations do not. Moreover, $\langle F\rangle=0$ by symmetry. As a result, the PMF is reduced to the external potential, $\nabla V_{*}=\nabla V_{ex}$, $E_{0}(t)$ coincides with $F_{0}(t)$, and the correlation $\langle F\,E_{0}(t)\rangle=\langle F\,F_{0}(t)\rangle$ does not depend on $X$. Clearly, that correlation also does not depend on $P$ and bath variables. Therefore, the propagator $e^{Lt}$ in the integral term of Eq. (81) can be dropped, and the equation takes the form of the standard generalized Langevin equation
$\dot{P}(t)=-\nabla V_{ex}[X(t)]+F_{0}(t)-\frac{\beta}{M}\int_{0}^{t}d\tau\,P(t-\tau)\,\langle F\,F_{0}(\tau)\rangle,$ (83)
supplemented with the ordinary fluctuation-dissipation relation.
For an anchored bath, the situation is more complicated because the statistical properties of noise $E_{0}(t)$ generally depend on $X$. While forces $F$ and $F_{0}(t)$ still depend on $X$ only through $X-x_{i}$, the internal bath forces depend not only on coordinate differences $x_{i}-x_{j}$, but also on the positions of individual bath units $x_{i}$. In that case, the variable change $x_{i}^{\prime}=x_{i}-X$ does not eliminate $X$ from the averages. As a result, $\langle F\rangle$ and $\langle FF_{0}(t)\rangle$ depend on $X$ and so does the correlation
$\langle FE_{0}(t)\rangle=\langle FF_{0}(t)\rangle-\langle F\rangle^{2}.$ (84)
One can reflect that dependence denoting
$\frac{\beta}{M}\langle FE_{0}(t)\rangle=\Gamma(t,X).$ (85)
Then Eq. (81) takes the form
$\dot{P}(t)=-\nabla V_{*}[X(t)]-E_{0}(t)(t)-\int_{0}^{t}d\tau\,P(t-\tau)\,\Gamma[\tau,X(t-\tau)],$ (86)
which is formally similar to the standard generalized Langevin equation (with the external potential replaced by the PMF), except that the dissipation term is not of the conventional convolution form since it depends on both $\tau$ and $t-\tau$. As it is, the equation is not operational because the dependence of the kernel on $X$ is a priory unknown. That is not necessarily a problem. After all, the exact form of the PMF $V_{*}(X)$ is usually not known either; it is often defined empirically to fit experimental data. One may define the function $\Gamma(X)$ empirically as well. However, the difficulty is that the two functions $V_{*}(X)$ and $\Gamma(X)$ are clearly not independent, yet there seems to be no general relation describing their connection.
The situation is simplified when the Hamiltonian $H_{0}$ has a special structure (33), $H_{0}=h_{1}(x,X)+h_{2}(X)$, and therefore the force $F_{0}(t)$ on the fixed system has the form
$F_{0}(t)=e^{L_{0}t}\left(-\frac{\partial H_{0}}{\partial X}\right)=f_{1}(X,x,t)+f_{2}(X),$ (87)
where component $f_{2}(X)$ depends only on the system’s position. In addition, let us assume that component $f_{1}$ has the properties of noise in the standard theory of Brownian motion. Namely, the average of $f_{1}$ over bath variables (with distribution $\rho_{0}$) is zero for any $t$,
$\langle f_{1}(X,x,t)\rangle=0,$ (88)
and its auto-correlation correlation
$\langle f_{1}(X,x,0)\,f_{1}(X,x,t)\rangle\equiv\langle f_{1}f_{1}(t)\rangle=c(t)$ (89)
does not depend on the system’s position $X$. In that case,
$\langle F\rangle=f_{2}(X),\quad E_{0}(t)=F_{0}(t)-\langle F\rangle=f_{1}(X,x,t),$ (90)
and the correlation $\langle F\,E_{0}(t)\rangle$ coincides with the auto-correlation of $f_{1}$
$\langle FE_{0}(t)\rangle=\langle(f_{1}+f_{2})\,f_{1}(t)\rangle=\langle f_{1}f_{1}(t)\rangle+f_{2}\langle f_{1}(t)\rangle=\langle f_{1}f_{1}(t)\rangle,$ (91)
and therefore, because of assumption (89), does not depend on $X$. Then in Eq. (80) the propagator $e^{L(t-\tau)}$ has no effect and can be dropped. Also, since noise $E_{0}(t)$ is zero-centered, the correlation $\langle FE_{0}(t)\rangle$ coincides with the auto-correlation of $E_{0}(t)$,
$\langle E_{0}(0)\,E_{0}(t)\rangle=\langle F\,E_{0}(t)\rangle-\langle F\rangle\,\langle E_{0}(t)\rangle=\langle F\,E_{0}(t)\rangle.$ (92)
As a result, Eq. (80), written in terms of the true momentum $P=P_{*}/\lambda$, turns into the standard generalized Langevin equation
$\dot{P}(t)=-\nabla V_{ex}[X(t)]+E_{0}(t)-\frac{\beta}{M}\int_{0}^{t}d\tau\,P(t-\tau)\,\langle E_{0}(0)\,E_{0}(\tau)\rangle$ (93)
with the external potential $V_{ex}(X)$ replaced by the PMF $V_{*}(X)$. This is the equation advocated in Refs. Lange ; Kinjo . We observe that the equation is valid only under special assumptions. In the next two sections we consider a special (linear) version of the model with an anchored bath where assumptions (87)-(89) do hold and therefore the Langevin equation (93) is valid.
VI Klein-Gordon chain: the PMF
Let us now consider a more specific model where the bath is formed by the Klein-Gordon harmonic chain. The chain consists of $N\gg 1$ atoms of mass $m$, connected by springs with stiffness $k_{0}$. Each atom of the chain is subjected to the on-site anchor harmonic potential with stiffness $k_{a}$. The system of interest is a particle of mass $M$, connected by a $k_{0}$-spring to the first atom of the chain and subjected to external potential $V_{ex}(X)$. We assume a boundary condition when the terminal atom of the chain is connected by a $k_{0}$-spring to a wall. The model is sketched in Fig. 1 and can be viewed as the familiar Rubin model Zwanzig ; Weiss , extended by the presence of anchor potentials acting on atoms of the bath and external potential $V_{ex}(X)$ acting on the system. It is well known that a non-anchored chain can serve, in the limit $N\to\infty$, as a passive bath Zwanzig ; Weiss . In contrast, the Klein-Gordon chain is non-passive: in thermal equilibrium it produces a systematic harmonic potential acting on the system, so the PMF has the form
$V_{*}(X)=V_{ex}(X)+\frac{1}{2}\,\kappa\,X^{2}.$ (94)
Below we evaluate stiffness $\kappa$ of the bath-induced potential; the result is given by Eq. (110).
We write the total Hamiltonian of the model in form (40),
$H=H_{s}+H_{0}(X),$ (95)
where $H_{s}$ is the Hamiltonian of the isolated system
$H_{s}=\frac{P^{2}}{2M}+V_{ex}(X),$ (96)
and $H_{0}(X)$ is the Hamiltonian of the bath interacting with the system fixed at position $X$,
$H_{0}(X)=\sum_{i=1}^{N}\left\{\frac{p_{i}^{2}}{2m}+\frac{k_{0}}{2}(x_{i}-x_{i+1})^{2}+\frac{k_{a}}{2}\,x_{i}^{2}\right\}+\frac{k_{0}}{2}\,(X-x_{1})^{2}.$ (97)
Coordinates $x_{i}$ represent displacements of the chain’s atoms from positions of mechanical equilibrium. In accordance with the chosen boundary condition, we assume in Eq. (97) that $x_{N+1}=0$. It is convenient to present $H_{0}$ as
$H_{0}=H_{0}^{\prime}+H_{0}^{\prime\prime},$
$H_{0}^{\prime}=\sum_{i=0}^{N}\left\{\frac{p_{i}^{2}}{2m}+\frac{k_{0}}{2}(x_{i}-x_{i+1})^{2}+\frac{k_{a}}{2}\,x_{i}^{2}\right\},\qquad H_{0}^{\prime\prime}=-k_{0}\,x_{1}\,X+\frac{k_{0}}{2}\,X^{2},$ (98)
assuming that $x_{0}=x_{N+1}=0$ and $p_{0}=0$. Hamiltonian $H_{0}^{\prime}$ describes the isolated Klein-Gordon chain of $N$ atoms with terminal atoms $i=1$ and $i=N$ connected to the walls, see Fig. 2.
Figure 1: The setup considered in Sections 6 and 7. The black circle represents the system of interest, which is a classical particle subjected to an (arbitrary) external potential $V_{ex}(X)$. The system is connected by the spring with stiffness $k_{0}$ to the thermal bath, formed by the Klein-Gordon lattice. The latter consists of $N$ atoms connected by the springs with stiffness $k_{0}$. Each atom of the lattice is subjected to on-site (anchoring) harmonic potentials with stiffness $k_{a}$.
The advantage of working with $H_{0}^{\prime}$ is that it can be readily diagonalized using the well-known normal mode transformation $\{x_{i},p_{i}\}\to\{Q_{j},P_{j}\}$,
$x_{i}=\frac{1}{\sqrt{m}}\,\sum_{j=1}^{N}A_{ij}\,Q_{j},\quad p_{i}=\sqrt{m}\,\sum_{j=1}^{N}A_{ij}\,P_{j},$ (99)
where the transformation matrix
$A_{ij}=\sqrt{\frac{2}{N+1}}\,\sin\frac{ij\pi}{N+1}$ (100)
satisfies the orthogonality relations
$\sum_{i=1}^{N}A_{ij}A_{ij^{\prime}}=\delta_{jj^{\prime}},\quad\sum_{j=1}^{N}A_{ij}A_{i^{\prime}j}=\delta_{ii^{\prime}}.$ (101)
The transformation has the same form as for the corresponding harmonic chain with no anchor potentials. It is canonical and preserves phase-space volumes, i. e. the Jacobian of the transformation equals $1$. Therefore, the average $\langle\cdots\rangle$ over the bath variables $\{x_{i},p_{i}\}$, which we used in the previous sections, can be replaced by the average over $\{Q_{j},P_{j}\}$.
Expressed in terms of $\{Q_{j},P_{j}\}$, Hamiltonians $H_{0}^{\prime}$ and $H_{0}^{\prime\prime}$ take the forms
$H_{0}^{\prime}=\frac{1}{2}\sum_{j=1}^{N}\left\{P_{j}^{2}+\Omega_{j}^{2}\,Q_{j}^{2}\right\},\qquad H_{0}^{\prime\prime}=-\sum_{j=1}^{N}c_{j}\,Q_{j}\,X+\frac{k_{0}}{2}\,X^{2},$ (102)
where the coupling coefficients $c_{j}$ are
$c_{j}=\frac{k_{0}}{\sqrt{m}}A_{1j}=\frac{k_{0}}{\sqrt{m}}\sqrt{\frac{2}{N+1}}\,\sin\frac{\pi j}{N+1},$ (103)
and the normal mode frequencies $\Omega_{j}$ are determined by the following expressions
$\Omega_{j}^{2}$ $=$ $\omega_{j}^{2}+\omega_{a}^{2},\quad\omega_{j}=\omega_{0}\sin\frac{\pi j}{2(N+1)},\quad\omega_{0}^{2}=4\,k_{0}/m,\quad\omega_{a}^{2}=k_{a}/m.$ (104)
In these expressions, $\omega_{j}$ are normal mode frequencies and $\omega_{0}$ is the highest normal mode frequency in the unanchored chain (with $\omega_{a}=0$). Note that for the anchored chain the spectrum of normal modes is bounded not only from above (as for the non-anchored chain), but also from below
$\omega_{a}^{2}<\Omega_{j}^{2}<\omega_{0}^{2}+\omega_{a}^{2}.$ (105)
This feature may be important for ergodic properties of the combined system Plyukhin_NEO .
Figure 2: The Klein-Gordon lattice described by Hamiltonian $H_{0}^{\prime}$ in Eq. (98).
With $H_{0}^{\prime}$ and $H_{0}^{\prime\prime}$ given by Eq. (102), Hamiltonian $H_{0}=H_{0}^{\prime}+H_{0}^{\prime\prime}$ takes the form
$H_{0}=\frac{1}{2}\sum_{j=1}^{N}\left\{P_{j}^{2}+\Omega_{j}^{2}\left[Q_{j}^{2}-\frac{2c_{j}}{\Omega_{j}^{2}}\,Q_{j}\,X\right]\right\}+\frac{k_{0}}{2}\,X^{2}.$ (106)
Completing the square in $\cdots$], one gets
$H_{0}=\frac{1}{2}\sum_{j=1}^{N}\left\{P_{j}^{2}+\Omega_{j}^{2}\left[Q_{j}-\frac{c_{j}}{\Omega_{j}^{2}}\,X\right]^{2}\right\}+\frac{1}{2}\,\left[k_{0}-\sum_{j=1}^{N}\left(\frac{c_{j}}{\Omega_{j}}\right)^{2}\right]X^{2}.$ (107)
In the limit $N\to\infty$ one finds
$\sum_{j=1}^{N}\left(\frac{c_{j}}{\Omega_{j}}\right)^{2}\to\frac{k_{0}}{\pi}\,\int_{0}^{\pi/2}\frac{\sin^{2}(2x)\,dx}{a^{2}+\sin^{2}x}=k_{0}(1+2a^{2}-2a\sqrt{1+a^{2}}),$ (108)
where we introduced the parameter $a=\omega_{a}/\omega_{0}$, characterizing the relative strength of the anchor potential. Then, in the limit of the infinite chain, $H_{0}$ takes the form
$H_{0}=\frac{1}{2}\sum_{j=1}^{N}\left\{P_{j}^{2}+\Omega_{j}^{2}\left[Q_{j}-\frac{c_{j}}{\Omega_{j}^{2}}\,X\right]^{2}\right\}+\frac{1}{2}\,\kappa\,X^{2},$ (109)
with
$\kappa=2\,k_{0}\,a\left\{\sqrt{1+a^{2}}-a\right\},\quad a=\frac{\omega_{a}}{\omega_{0}}=\sqrt{\frac{k_{a}}{4\,k_{0}}}.$ (110)
We observe that $H_{0}$ has the form (33), discussed at the end of Section 2, with positions of the bath particles $x=\{x_{i}\}$ replaced by normal modes $Q=\{Q_{j}\}$,
$H_{0}=h_{1}(Q,X)+h_{2}(X),$
$h_{1}=\frac{1}{2}\sum_{j}\left\{P_{j}^{2}+\Omega_{j}^{2}\left[Q_{j}-\frac{c_{j}}{\Omega_{j}^{2}}\,X\right]^{2}\right\},\qquad h_{2}(X)=\frac{1}{2}\,\kappa\,X^{2}.$ (111)
Using the substitution $Q^{\prime}_{j}=Q_{j}-\frac{c_{j}}{\Omega_{j}^{2}}\,X$, one observes that the partition function corresponding to $h_{1}$
$Z_{1}=\operatorname{Tr}_{b}\left(e^{-\beta h_{1}}\right)$ (112)
does not depend on $X$. Then the partition function $Z_{0}$ corresponding to $H_{0}$ has the form (35),
$Z_{0}(X)=\operatorname{Tr}_{b}\left(e^{-\beta H_{0}}\right)=Z_{1}\,e^{-\beta\,h_{2}(X)},$ (113)
Defining the constant $c$ in the expression (13) of the bath-induced potential $\Delta V$ to be $c=Z_{1}$, we find that
$\Delta V=-\frac{1}{\beta}\ln\left[\frac{Z_{0}(X)}{c}\right]=h_{2}(X)=\frac{1}{2}\,\kappa\,X^{2},$ (114)
and the PMF $V_{*}=V_{ex}+\Delta V$ has the form (94).
Thus, the bath formed by the infinite Klein-Gordon chain is non-passive; it acts on the system with the systematic average force $\langle F\rangle=-\kappa\,X$. The effective stiffness $\kappa$ of that force as a function of $a=\omega_{a}/\omega_{0}$ is given by Eq. (110). It monotonically increases from 0 at $a=0$ to $k_{0}$ for $a\gg 1$. Both limiting values are intuitively plausible. For $a=0$, the model is equivalent to the Rubin model Zwanzig ; Weiss where the chain is not anchored and in the limit $N\gg 1$ (but not for finite $N$) serves as a passive bath, which does not produce a systematic force on the fixed system. For $a\gg 1$ the chain acts as an infinitely heavy anchor. In that case, the system feels only the spring of stiffness $k_{0}$ by which it is attached to the terminal atom of the chain.
VII Klein-Gordon chain: Langevin equation
We have seen at the end of Section 5 that a single-particle system subjected to a PMF is described (in the leading order in $\lambda$) by the standard generalized Langevin equation (93) if statistical properties of noise
$E_{0}(t)=F_{0}(t)-\langle F\rangle,$ (115)
specifically the correlation $\langle E_{0}(0)\,E_{0}(t)\rangle$, do not depend on the system’s variables. Let us show that this is just the case for the model described in the previous section, i.e. when the bath, linearly coupled to the system, is formed by the harmonic Klein-Gordon chain.
The microscopic force exerted by the bath on the system $F=-k_{0}(X-x_{1})$ can be written in terms of normal modes as follows
$F=-\frac{\partial H_{0}}{\partial X}=-\frac{\partial H_{0}^{\prime\prime}}{\partial X}=-k_{0}X+\sum_{j}c_{j}Q_{j},$ (116)
where $c_{j}$ are given by Eq. (103). Then $F_{0}(t)=e^{L_{0}t}F$ has the form
$F_{0}(t)=-k_{0}X+\sum_{j}c_{j}Q_{j}(t).$ (117)
The equations of motion for $Q_{j}(t)$, governed by Hamiltonian $H_{0}$, Eq. (106), read
$\ddot{Q}_{j}(t)+\Omega_{j}^{2}Q_{j}(t)=c_{j}X,$ (118)
where the system’s position $X$ is a fixed parameter. Solutions of these equations can be written as
$Q_{j}(t)=\left(Q_{j}-\frac{c_{j}}{\Omega_{j}^{2}}\,X\right)\,\cos\Omega_{j}t+\frac{1}{\Omega_{j}}P_{j}\,\sin\Omega_{j}t+\frac{c_{j}}{\Omega_{j}^{2}}X,$ (119)
where $Q_{j}=Q_{j}(0)$ and $P_{j}=\dot{Q}_{j}(0)$. Substituting to Eq. (117) yields
$F_{0}(t)=\sum_{j}c_{j}\left\{\left(Q_{j}-\frac{c_{j}}{\Omega_{j}^{2}}\,X\right)\,\cos\Omega_{j}t+\frac{1}{\Omega_{j}}P_{j}\,\sin\Omega_{j}t\right\}-\kappa X$ (120)
with
$\kappa=k_{0}-\sum_{j=1}^{N}\left(\frac{c_{j}}{\Omega_{j}}\right)^{2}.$ (121)
As was shown in the previous section, in the limit of the infinite chain, $N\to\infty$, the effective stiffness $\kappa$ takes the form (110).
We observe that $F_{0}(t)$ satisfies conditions (87)-(89), mentioned at the end of Section 5, which ensure that statistical properties of the zero-centered noise $E_{0}(t)=F_{0}(t)-\langle F\rangle$ do not depend on $X$. We can also show that directly. Indeed, using the substitution
$Q_{j}^{\prime}=Q_{j}-\frac{c_{j}}{\Omega_{j}^{2}}\,X,\quad P_{j}^{\prime}=P_{j},$ (122)
we observe that in terms of $Q_{j}^{\prime},P_{j}^{\prime}$ the Hamiltonian $H_{0}$ has the quadratic form
$H_{0}=\frac{1}{2}\sum_{j}\left\{{P_{j}^{\prime}}^{2}+\Omega_{j}^{2}\,Q_{j}^{\prime 2}\right\}+\frac{1}{2}\,\kappa\,X^{2},$ (123)
so that
$\langle Q^{\prime}_{j}\rangle=0,\quad\langle Q^{\prime}_{j}Q^{\prime}_{k}\rangle=\frac{\delta_{jk}}{\beta\,\Omega_{j}^{2}},\quad\langle P^{\prime}_{j}P^{\prime}_{k}\rangle=\frac{\delta_{jk}}{\beta}.$ (124)
From these relations and Eq. (120) one obtains
$\langle F\rangle=-\kappa X.$ (125)
Therefore, the zero-centered noise $E_{0}(t)=F_{0}(t)-\langle F\rangle$ has the form
$E_{0}(t)=\sum_{j}c_{j}\left\{Q^{\prime}_{j}\,\cos\Omega_{j}t+\Omega_{j}^{-1}P_{j}^{\prime}\,\sin\Omega_{j}t\right\},$ (126)
and its correlation
$\langle E\,E_{0}(t)\rangle=\frac{1}{\beta}\,\sum_{j=1}^{N}\left(\frac{c_{j}}{\Omega_{j}}\right)^{2}\cos\Omega_{j}t$ (127)
does not depend on $X$. In that case, as discussed in Section 5, the Langevin equation takes the standard form (93), which we can write as
$\dot{P}(t)=-\nabla V_{*}[X(t)]+E_{0}(t)-\int_{0}^{t}d\tau\,P(t-\tau)\,\Gamma(\tau),$ (128)
where the PMF equals
$V_{*}(X)=\frac{1}{2}\,\kappa X^{2}+V_{ex}(X),$ (129)
stiffness $\kappa$ of the bath-induced potential is given by Eq. (110), and the dissipative kernel $\Gamma(t)$ is related to the zero-centered noise $E_{0}(t)$ by the standard fluctuation-dissipation relation
$\Gamma(t)=\frac{\beta}{M}\,\langle EE_{0}(t)\rangle.$ (130)
Recall that we derived the Langevin equation (128) expanding the projected force in powers of $\lambda=\sqrt{m/M}$
$F^{\dagger}(t)=e^{\mathcal{Q}Lt}F=E_{0}(t)+\lambda E_{1}(t)+\cdots$ (131)
and restricting to the leading approximation $F^{\dagger}(t)\approx E_{0}(t)$. One can show that for the Klein-Gordon model the next term $E_{1}(t)$, which is given by Eq. (78), vanishes. Indeed, as follows from Eqs. (122) and (126), $E_{0}(t)$ is linear in $X$, the derivative
$\frac{\partial E_{0}(t)}{\partial X}=\sum_{j}c_{j}\,\frac{\partial Q_{j}^{\prime}}{\partial X}=-\sum_{j}\left(\frac{c_{j}}{\Omega_{j}}\right)^{2}$ (132)
does not depend on bath variables, and therefore
$E_{1}(t)\sim\mathcal{Q}\,\ \frac{\partial E_{0}(t)}{\partial X}=0.$ (133)
In a similar way one can show that all terms of higher order in the expansion (131) vanish as well. Thus, for the presented model the projected force equals $E_{0}(t)$ exactly,
$F^{\dagger}(t)=E_{0}(t),$ (134)
and the Langevin equation (128) is also exact.
Figure 3: The dimensionless dissipation kernel $\gamma(t)$, see Eq. (136), for three values of the anchor parameter $a=\omega_{a}/\omega_{0}$.
The derivation produced above is, of course, not the most economical. We used the projection operator technique with the hope (unfulfilled) to get a Langevin equation for a general setting described in Section 3. Had we assumed the linearity of forces from the start, we could have derived the Langevin equation (128) more easily by the explicit integration method, i. e. by integrating linear equation of motion of the bath Zwanzig ; Weiss ; Hanggi . We address that method in the Appendix.
We finish this section by deriving a more explicit expression for the correlation $\langle EE_{0}(t)\rangle$ and the dissipation kernel $\Gamma(t)$. According to Eqs. (127) and (130), the kernel equals
$\Gamma(t)=\frac{1}{M}\,\sum_{j=1}^{N}\left(\frac{c_{j}}{\Omega_{j}}\right)^{2}\cos\Omega_{j}t.$ (135)
Taking into account Eqs. (103) and (104) for $c_{j}$ and $\Omega_{j}$, and taking the limit $N\to\infty$ we obtain
$\Gamma(t)=\lambda^{2}\omega_{0}^{2}\,\gamma(t),\qquad\gamma(t)=\frac{1}{4\pi}\,\int_{0}^{\pi/2}\frac{\sin^{2}(2x)}{a^{2}+\sin^{2}x}\,\cos\left(\omega_{0}t\sqrt{a^{2}+\sin^{2}x}\right)\,dx,$ (136)
and the noise correlation equals $\langle EE_{0}(t)\rangle=\frac{M}{\beta}\,\Gamma(t)$. For the non-anchored chain ($a=0$), the dimensionless kernel $\gamma(t)$ is expressed in terms of the Bessel function $J_{1}(x)$,
$\gamma(t)=\frac{1}{2\,\omega_{0}\,t}\,J_{1}(\omega_{0}t).$ (137)
This is the well-known result for the Rubin model Zwanzig ; Weiss . For nonzero $a$ the kernel (136) cannot be evaluated analytically. Fig. 3 presents numerical evaluations of $\gamma(t)$ for several values of $a$.
VIII Conclusion
The concept of the PMF naturally emerges in equilibrium statistical mechanics of complex systems consisting of many particles. The force $F_{i}$ on $i$-th particle of the system is produced by the bath disturbed by the interaction with other parts of the system. As a result, even if the bath is a translationally-invariant fluid, the mean force $\langle F_{i}\rangle$ is generally not zero and is determined by the bath-induced part of the PMF, denoted above as $\Delta V$. With the PMF it is straightforward to evaluate average values in equilibrium, but there is a variety of views on how to incorporate the PMF in the Langevin dynamics.
In this paper, we have simplified the problem by assuming that the system of interest consists of a single particle. In that case, the PMF is usually irrelevant because the homogeneous bath is passive and produces no systematic force on a fixed system. That is not so, however, if the bath is not translationally-invariant. In the presented model the symmetry of the bath is broken by on-site anchor potentials applied to individual particles of the bath. Although that setting offers a simpler exposition, we were unable to derive a Langevin equation in a closed form. The non-passivity of the bath not only modifies the average potential acting on the system, but also makes the dissipation kernel and noise statistics depend on the position of the system $X$. That dependence is a priory unknown and there seems to be no general relation connecting it to the PMF. Therefore, we argue that there is no general form for a Langevin equation with the PMF. The conclusion corroborates the findings of Refs. Schilling ; Schilling_review ; Vroy , although we used here a different projection operator technique.
An exception is the case where all internal forces in the combined system are linear. As a specific example, we considered the bath formed by the harmonic Klein-Gordon chain. In that case, the dissipation kernel and noise statistics do not depend on $X$, and the Langevin equation has the standard form with the external potential replaced by the PMF. Such an equation is often exploited in empirical studies, but we observe that it holds only under rather exceptional conditions.
Appendix
In Section 7 we derived the Langevin equation for the model where the bath is represented by the harmonic Klein-Gordon chain by applying the generic method based on the projection operator technique. In this Appendix we present an alternative method based on explicit integration of equations of motion of the bath. Of course, that method is only feasible when the equations of motion are linear. The derivation is similar to those for the Caldeira-Leggett and Rubin models Zwanzig ; Weiss ; Hanggi .
The Hamiltonian $H$ of the model is given by Eqs. (95)-(97), and the equation of motion for the system is
$\dot{P}=-\frac{\partial H}{\partial X}=-k_{0}\,(X-x_{1})-\frac{\partial V_{ex}}{\partial X}.$ (A1)
Using transformation (99) to the bath’s normal modes, the equation can be written as
$\dot{P}=-k_{0}\,X-\sum_{j}c_{j}Q_{j}-\frac{\partial V_{ex}}{\partial X},$ (A2)
where $c_{j}$ are given by Eq. (103). The equations of motion for the bath’s normal modes are
$\dot{P}_{j}=-\frac{\partial H}{\partial Q_{j}}=-\Omega_{j}^{2}Q_{j}+c_{j}X,\quad\dot{Q}_{j}=\frac{\partial H}{\partial P_{j}}=P_{j},$ (A3)
Excluding $P_{j}$, one obtains
$\ddot{Q}_{j}(t)+\Omega_{j}^{2}\,Q_{j}(t)=c_{j}X(t).$ (A4)
The general solution has the form
$Q_{j}(t)=Q_{j}^{0}(t)+\frac{c_{j}}{\Omega_{j}}\,\int_{0}^{t}\sin(\Omega_{j}\tau)\,X(t-\tau)\,d\tau,$ (A5)
where $Q_{j}^{0}(t)$ is the general solution of the homogeneous equation $\ddot{Q}_{j}+\Omega_{j}^{2}\,Q_{j}=0$,
$Q_{j}^{0}(t)=Q_{j}(0)\cos(\Omega_{j}t)+\frac{P_{j}(0)}{\Omega_{j}}\,\sin(\Omega_{j}t).$ (A6)
Integrating by parts, Eq. (A5) can be written as
$Q_{j}(t)=Q_{j}^{0}(t)+\frac{c_{j}}{\Omega_{j}^{2}}\left\{X(t)-X(0)\,\cos(\Omega_{j}t)-\int_{0}^{t}\cos(\Omega_{j}\tau)\,\dot{X}(t-\tau)\,d\tau\right\}.$ (A7)
Substituting into Eq. (A2) yields
$\dot{P}(t)=-\frac{\partial V_{ex}}{\partial X}-\kappa\,X(t)-\int_{0}^{t}\Gamma(t-\tau)\,P(\tau)\,d\tau+\zeta(t)-M\,X\,\Gamma(t),$ (A8)
where the dissipation kernel $\Gamma(t)$ and noise $\zeta(t)$ are
$\Gamma(t)=\frac{1}{M}\,\sum_{j=1}^{N}\left(\frac{c_{j}}{\Omega_{j}}\right)^{2}\cos(\Omega_{j}t),\quad\zeta(t)=\sum_{j=1}^{N}c_{j}Q_{j}^{0}(t),$ (A9)
the effective stiffness $\kappa$ is
$\kappa=k_{0}-\sum_{j=1}^{N}\left(\frac{c_{j}}{\Omega_{j}}\right)^{2},$ (A10)
and $X=X(0)$. In the limit $N\to\infty$ $\kappa$ is given by expression (110). Finally, absorbing the term $M\,X\,\Gamma(t)$ into noise and using notations of the main text, we obtain
$M\ddot{X}(t)=-\nabla V_{*}[X(t)]-\int_{0}^{t}\Gamma(t-\tau)\,P(\tau)\,d\tau+E_{0}(t),$ (A11)
with the PMF
$V^{*}=V_{ex}+\frac{1}{2}\kappa q^{2},$ (A12)
and noise
$E_{0}(t)=\zeta(t)-M\,X\,\Gamma(t)=\sum_{j=1}^{N}c_{j}\,\left[\left(Q_{j}-\frac{c_{j}}{\Omega_{j}^{2}}\,X\right)\,\cos(\Omega_{j}t)+\frac{P_{j}}{\Omega_{j}}\,\sin(\Omega_{j}t)\right],$ (A13)
where $X=X(0)$, $Q_{j}=Q_{j}(0)$, $P_{j}=P_{j}(0)$. Eq. (A11) coincides with the Langevin equation (128).